CuSO4 + Fe = ?

CuSO4 + Fe = ?

Grade:10

4 Answers

KUNCHAM SAMPATH
8 years ago
Fe2(So4)3(aq)CuSO4 + Fe =
KUNCHAM SAMPATH
8 years ago
sorry.....CuSO4 + Fe = Fe2(So4)3(aq)
RAKESH CHINDAM
8 years ago
CuSO4+Fe = Fe2(SO4)3(aq) 
SHANMUKESHWAR
8 years ago
 CuSO4+Fe = Fe2(SO4)3(aq)